# An engineering guide to superconducting quantum circuit shielding

### v1

@@ II.1.3 Practical implementation of IR shielding
We considered seven shielding configurations with the increased complexity (weight and cost). Shielding configuration "A" consists solely of the sample holder and the can on the lower stage. This represents the simplest possible configuration. Shielding system "B" incorporates an absorbing coating on the can the can. Comparison with system "A" enables evaluation of thermal radiation absorption effectiveness of the lower cryostat stage shield. Configurations "C"-"E" examine different absorbing coating placements: "C" – Absorbing coating on the inner surface of sample holder; "D" - outer surface of sample holder; "E" - both surfaces. These configurations, compared with "A", determine optimal coating placement on the sample holder. Configuration "F" adds a shielded enclosure with inner absorbing coating around the optimal sample holder configuration (selected from "A", "C"-"E"). This tests the effectiveness of the shield with absorbing coating surrounding the holder. Configuration "G" introduces a second shielded layer with absorbing coating. Comparison with F assesses the benefit of additional shielding layers.
The theoretical calculation and modeling results for the shielding systems with various absorbing coating positions are represented in Fig. 6 and demonstrate similar results with a mean difference of approximately 6.1%. The sample temperature is represented in relative units in order to demonstrate the its temperature dependence on different shielding configurations relative to the worst-case shielding.
Based on the results, it can be seen that the nearest surface to the sample has a significant influence on its temperature. In our case, it is the sample holder lid, that is why it should have an absorbing surface. Starting from the “E” shielding configuration, it can be seen that the sample temperature is not changing (for shielding configurations “E”, “F”, and “G”), i.e., a further increasing of IR shielding quantity does not lead to an improvement in the shielding effectiveness. Therefore, an additional IR shields in the shielding system are unnecessary, as the IR protection provided by the sample holder lid appears to be sufficient. Another important conclusion that follows from the above discussion is that the sample holder itself is also a shielding system component, and should therefore be considered along with the shields.
Figure 6: Calculated sample temperature for different shielding configurations (simulations and theoretical results).
According to our review, for a high effectiveness IR shields must be light-tight Hover (2013). At least in a half of considered papers, where the shield’s shape is indicated, they consist of a cup with a lid Place et al. (2021); Zhang (2018). In some studies, there are special mentions, where the IR shield is hermetically sealed Geerlings (2013), light-tight Chow et al. (2014); Lane et al. (2020); Megrant et al. (2012) or the connection between the shield and the lid is sealed by the indium wire Geerlings (2013); Harrington (2020). Much less frequently, cylinder-shaped inserts are used Pop et al. (2014); Geerlings (2013), in rare cases, an individual shield may be used above the SQC Arute et al. (2019) or around it in the form of a tube Chou (2018); Wang et al. (2019); Serniak (2019). Technically, in order to ensure a light-tightness (or “IR-radiation-tight”), IR shields can be effectively integrated into standard dilution refrigerators, when they are designed as a cylinder with a lid. The connection between the cylinder shields and the lid can be threaded and easily sealed using the indium wire, and control wiring can be dropped into the shielded space through hermetic connectors in the lid. From the one hand, it is important to avoid any gaps (comparable to IR wavelength) between the lid and shields to ensure IR-radiation-tight solution. From the other hand, one should think on special form designed path for pumping down the shielded space.
@@
In order to ensure a robust thermal contact, approximately in a half of considered papers, oxygen-free copper Zaretskey et al. (2013); Buks et al. (2020); Scigliuzzo et al. (2020); Eddins et al. (2019); Underwood (2015) is used (see Fig. 7). The choice of a copper is justified by its high thermal conductivity. The holders manufactured of superconductors, such as aluminum Jin et al. (2015); Lane et al. (2020); Cai et al. (2019); Burkhart et al. (2021); Lucero et al. (2012), considered in the other half of considered papers. The use of superconductors creates additional magnetic shielding owing to the Meissner effect (more detail consideration of this effect will be done later). Apart from aluminum, niobium Srinivasan et al. (2014) or titanium alloy Kumar et al. (2016) can also be used as superconductors, although such instances are rare. Despite the added protection from the magnetic field, superconducting sample holders have a lower thermal conductivity compared to the copper sample holders, that is why the use of the superconductors is undesirable. There are also copper sample holders coated by superconductor (lead Anton et al. (2012) or indium Narla (2017)). These sample holders create additional magnetic shielding while maintaining the high thermal conductivity. However, these examples are not common. There are also a few cases, where a copper coated with gold is used Yamamoto et al. (2014); Lienhard et al. (2019); Blok et al. (2021). The gold coating prevents the formation of the oxides on the copper, thereby reducing dielectric losses Huang et al. (2021).
@@
Among the works, where the shape of the sample holder is shown, approximately in a half of them the transmon qubits housed in a three-dimensional cavity Moghaddam et al. (2019); Martin et al. (2020); Monroe et al. (2021b); Merkel et al. (2013); Serniak et al. (2018), which serves as both a chip holder and a qubit circuit element for readout. The remaining works have three types of sample holder shapes: approximately 20% use the rectangular holders de Visser (2014); Yeh et al. (2017); Ribeill (2016), 16% use the cylindrical holders Chow (2010); Chen (2018); Reed (2013), and 11% use a flat copper bases with printed circuit boards (PCBs) and lids above the chip Place et al. (2021); Ann (2021); Huang et al. (2021).
The selection of a sample holder’s shape depends on the SQC design and the wiring method to the cryogenic setup. Only the quality of the sample holder internal surface and its sealing are of importance.
In considered articles, just 6% of the sample holders have the lid with an IR absorbent coating. This solution is in good agreement with our calculations performed earlier. For this purpose, Stycast epoxy with additional SiC particles de Visser (2014), a mixture of SiC and finely dispersed carbon powder Barends et al. (2011), or Eccosorb CR-124 electromagnetic absorber Córcoles et al. (2011); Srinivasan (2013) can be used. The absorbing coating on the inner surface of the sample holder allows for more efficient cooling of the SQC, and at the same time it is also important that the holder has a high thermal conductivity. However, high dielectric loss of the absorber can affect qubit‘s coherence if they are close to each other. IR absorber should be positioned sufficiently far from the qubit. Sufficient distance can be determined using surface participation method Lienhard et al. (2019); Smirnov et al. (2024). This approach shows that the minimal qubit-absorber distance should exceed 4 mm. The surface participation method allows to quantify the extent to which the absorber interacts with the qubit’s field in terms of maximal possible coherence time T1 that can be achieved in presence of the dielectric loss channel:
@@
where $T_{1}$, $\omega$, and $Q$ are coherence time, angular frequency and quality factor, respectively. $tan{\delta}_{abs}$ is the dielectric loss tangent of the absorber, $p_{abs}$ is a participation ratio of the material (a fraction of electromagnetic field energy stored within this material). ${\Gamma}_{0}$ represents other loss channels. The simulation conducted with a typically sized floating transmon Smirnov et al. (2024) and a usual transmon on a lossless substrate in a lossless chip cavity to ensure ${\Gamma}_{0}=0$. Inner side of the top lid was covered with RF absorber with thickness of 0.5 mm and $tan{\delta}_{abs}$ of 0.05, which is an upper bound for STYCAST IR absorber. During the simulation distance between the qubit and the bottom surface of the absorber was varied from 0.5 to 5 $mm$. The modelling shows that the influence of the absorber becomes as low as an influence of dielectric losses in silicon (limiting $T_{1}$ more than 10 $ms$) substrate when STYCAST is more than 4 $mm$ away from the qubit (see Fig. 8). This ensures that the benefits of efficient cooling provided by the absorber are not compromised by degradation in qubit coherence.
@@
In some cases, a weakly magnetic Eccosorb foam Uilhoorn (2021) may be used wrapped around the holder to absorb electromagnetic waves. However, without a base with a high thermal conductivity this foam may not be cooled sufficiently leading to re-emission of absorbed radiation into the SQC. In 3% of the works reviewed, a seal is used Smith (2019); de Visser (2014); Lisenfeld et al. (2019) (particularly, indium wire Sears (2013)) to ensure light tightness.
Figure 9: Sample temperature versus gap size for different lid implementations: a) different holder designs; b) simulation results.
The connection of the sample holder and the lid is critical. In order to determine the optimal method of connecting the sample holder to the lid, we have modelled and analyzed their different connection implementations. Three variants of the lid for the cylindrical sample holder have been considered and corresponding models have been developed (see Fig. 9a).
The lid can either be common to the entire sample holder, including the chip, printed circuit board (PCB), and microwave connectors, or it can be individual covering only the sample with SQC. For the “standard lid”, the connector as a heat source (Heat source 1) is located inside the sample holder, and there is also an external heat source (Heat source 2), which can be the element of the measurement circuit, e.g. microwave filter. This model corresponds to the lid that covers the chip along with the PCB and connectors.
For “individual lid”, both heat sources are located outside the sample holder.
The “individual lid with overlap” has an additional overlap to prevent radiation from penetrating through the gap (see Fig. 9a). The gap between the lid and the sample is varied from 0 to 1 $mm$ in increments of 0.1 $mm$. As well as for an IR shielding modeling, relative units were used to represent the sample temperature.
Modeling results (Fig. 9b) demonstrate, that for a standard lid the sample temperature is approximately twice as high as compared to individual lid. The sample temperature increases gradually depending on the gap size. When choosing an individual lid, it is necessary to provide that there is no gap between the lid and the sample holder, because even 0.1 $mm$ gap elevates the sample temperature by 12% in the case of individual lid with overlap (or by 37% in the case of usual individual lid). These simulation results demonstrate the necessity of using an indium wire seal between the sample holder and the lid, that allows filling all small gaps and surface irregularities and so eliminate it. An individual lid with an overlap provides a lower sample temperature, especially about 26% lower than that at a standard individual lid.
To summarize, the sample holder for the SQC should be made of a copper (or a copper coated with aluminum), and the inner surface of the holder should be coated by Stycast epoxy mixed with 1 $mm$ SiC particles. Additionally, the sample holder lid of should be sealed with an indium wire to ensure light tightness.
@@ II.2.1 Shielding principle
Generally, protection from electrical and magnetic fields consists of three components: electrostatic shielding, magnetostatic shielding, and electromagnetic shielding (Fig. 10). The principle of shielding from an electrical field (Fig. 10a) involves the accumulation of the charge from a free space on the shield and its subsequent leakage into the ground. For effective protection against an electrical field, the shield must be made of a highly conductive material, such as copper or aluminum, with high-quality electrical grounding to minimize the contact resistance with the ground.
The shields made of ferromagnetic materials, such as permalloys (${\mu}$-metal, a soft magnetic alloy, or steels), with a high relative magnetic permeability (${\mu}_{r}$) are used to protect from a static and slowly varying (up to 1 $kHz$) magnetic field. In these shields, the lines of magnetic flux pass along the walls (see Fig. 10b), which have a lower magnetic resistance than the surrounding air. The shielding effectiveness is determined by magnetic permeability of the material the shield is made of.
The operation principle of the shields protecting from an alternating high-frequency magnetic field is based on an alternating electromagnetic field excitation within the shield. It creates alternating induction eddy currents (also known as Foucault currents). These currents shield the volume which this currents surround: the field inside the shield is brought down, while the field outside the shield is enhanced (see Fig. 10c). As a result, the field strength is reduced inside the shield and increased outside, thus leading to the field to be pushed from the shield. This type of shielding depends on the depth of field penetration at various frequencies (known as the "skin depth"), and becomes effective above frequencies ($f$) 1 $kHz$. The lower the skin depth ($\delta$) the thinner (the lighter) the shield can be made:
@@ II.2.3 Practical implementation of electromagnetic shielding
Superconducting shields used for protection from parasitic magnetic fields are primarily manufactured from aluminum (Fig. 11) Jin et al. (2015); Moghaddam et al. (2019); Asaad et al. (2016); Lanting et al. (2014); Sullivan et al. (2013); Liu (2016b). Other superconducting materials that were found in the literature: lead Burnett (2013); Anton et al. (2012); Jerger (2013), tin Kreikebaum et al. (2016) or their alloys de Visser (2014); Long (2019), niobium Murch et al. (2013) and indium Premkumar et al. (2021). Aluminum is preferred material for superconducting shielding due to the ease of its workability and relatively low cost compared to other materials.
In considered papers, almost a quarter of all superconducting shields in combination with an IR shielding, which are coated with an IR-absorbing material, are used (Fig. 12a) Zhang (2018); Arute et al. (2019); Kreikebaum et al. (2016). These shields usually consist of a light tight cup with a lid Place et al. (2021); Zhang (2018), which allows for more compact shielding due to the possibility of combining together IR and magnetic shielding (shielding from magnetic field). However, superconductors have a lower thermal conductivity than in its normal state. It has a negative impact on the absorbed heat dissipation, which is crucial for IR shielding. To effectively cool a superconducting shield, a superconducting layer can be placed on a copper base (Fig. 12b) Premkumar et al. (2021); Kreikebaum et al. (2016); Long (2019).
@@
Figure 12: Shielding systems with various IR shields: a) a 4-layer shielding with two IR-reflective Mylar layers and a combination of superconducting and IR shields Place et al. (2021); b) a 3-layer shielding: a superconducting shield combined with an IR shield and double-layered $\mu$-metal Premkumar et al. (2021); c) a single-layered cryoperm magnetic shield in combination with an IR absorber Córcoles et al. (2015).
@@
The operation principle of magnetic shielding is based on the magnetic field retraction into a shield material that has a high relative magnetic permeability (typically 100,000 or more) Huang et al. (2021). These materials are typically made of ferrum-nickel alloys that have been subjected to special processing, such as annealing in a hydrogen atmosphere. In the literature, these materials are often referred to as ${\mu}$-metals Burnett (2013); Hazard (2019); Yeh et al. (2017); Mirhosseini et al. (2020); Jr. (2018) or they may have specific names (Fig. 13), such as Cryoperm Lisenfeld et al. (2019); Sung (2018); Roch et al. (2014), Cryophy Dickel et al. (2018); Asaad et al. (2016); Dickel (2018), Amumetal Wang et al. (2019); Narla (2017); Liu (2016b); Shankar et al. (2013); Minev et al. (2019) (or Amuneal Córcoles et al. (2015); Chou (2018); Geerlings (2013)).
The presented materials differ in the method of their thermal processing in a hydrogen environment, and consequently, in their relative magnetic permeability (Table 2).
For these shields, it is preferable to use a material with the highest relative magnetic permeability. As an exception, an iron shield Lachance-Quirion et al. (2017) can be used. However, its relative magnetic permeability is much less compared to ${\mu}$-metals.
For magnetic shielding, cup-shaped shields are frequently used Rigetti et al. (2012); Bengtsson (2020). Cylindrical shields with a lid Underwood (2015); Kudra et al. (2020); Reagor (2015) are less frequently employed. Individual magnetic shielding is used in special cases Arute et al. (2019) As with superconducting shielding, the presence of a lid on the shield enables the equalization of the magnetic field within the protected space, thus it is recommended to use the magnetic shield with a lid.
@@
Iron (99.8% pure) 5,000 Le, Slodička, and Van Bockstal (2021)
Magnetic shields can also be combined with IR shielding (Fig. 12c) Córcoles et al. (2015); Beck (2018); Chow et al. (2014), but it was found only in 6% of considered papers. Such combination of shielding methods is not justified due to the low thermal conductivity of a ${\mu}$-metal compared to a copper, and can lead to a poor cooling of the absorbing coating. In some studies, specific guidelines are provided for the use of magnetic shields. For example, it is recommended not to include magnetic components inside the shield Kudra et al. (2020); Opremcak (2020); Gao (2018), or the sample holder should be placed closer to the bottom of the shield Underwood (2015), and a thermal anchor should be installed on the shield to improve its cooling Zhang (2018); Reagor (2015). Some magnetic shields have special design features such as waveguides on the lid Ann (2021); Dickel (2018) or light tightness Hazard (2019); Chow et al. (2014). These design features help to minimize the influence of the magnetic field on the SQC. Almost a half of considered shielding systems have the sample holders with special features. Either the holder itself is made of a superconductor Chen (2018); Mundhada (2020); Wilen et al. (2021), or it has superconducting Place et al. (2021); Anton et al. (2012) or absorbing layers Córcoles et al. (2011); Srinivasan (2013); de Visser (2014). These aspects have been discussed in more details before in a section about sample holders.
Figure 14: Simulation of magnetic field distribution within cylindrical shields: a) calculation model (the external magnetic field of 50 $\mu T$ directed along the axis of the cylinders); b) comparison of magnetic field strength along the shield axis for various shielding configurations; magnetic field distribution within: c) single $\mu$-metal shield (color bar screenshot from COMSOL for magnetic field strength is also valid for d)-g) configurations; the boundaries of the simulated shields are indicated by a black dashed line here and on the d)-g) various configurations: d) single aluminum shield, e) double $\mu$-metal shield, f) aluminum shield surrounded with a $\mu$-metal shield, g) $\mu$-metal shield surrounded with an aluminum shield.

### v2

@@ II.1.3 Practical implementation of IR shielding
We considered eight shielding configurations ordered by complexity increase (weight and cost) (Figure 5). Shielding configuration "A" consists solely of the sample holder and the can on the lower stage. This represents the simplest possible configuration. Shielding system "B" incorporates an absorbing coating on the can. Comparison with system "A" enables evaluation of thermal radiation absorption effectiveness of the lower cryostat stage shield. Configurations "C"-"E" examine different absorbing coating placements: "C" – Absorbing coating on the inner surface of sample holder; "D" - outer surface of sample holder; "E" - both surfaces. These configurations, compared with "A", determine optimal coating placement on the sample holder. Configuration "F" is the most frequently reported variant of IR shielding and it consists of the sample holder without absorbing coating surrounded by a shielded enclosure with an inner absorbing coating. Configuration "G" incorporates the same inner-coated enclosure, but applied to the optimal sample holder design (selected from configurations "A" and "C"-"E"). This arrangement allows evaluation of the effectiveness of the shield with absorbing coating surrounding the holder and comparison with widespread shielding system "F". Configuration "H" introduces a second shielded layer with absorbing coating. Configuration "H" in comparison with "G" assesses the benefit of additional shielding layers.
In both thermal scenarios sample temperature in the steady-state regime was evaluated using theoretical calculation and numerical modelling (Figure 6). The shielding system efficiency can be estimated by the minimum attainable temperature of the sample. In order to consider only radiative heat transfer the sample was deliberately decoupled from the refrigerator lowest plate by omitting thermal conduction paths. As a result, in each scenario sample temperature was greater than 260 $mK$ that significantly exceeds 10 $mK$ on the refrigirator’s lowest stage and is not usually observed in experiments. However, this approach provides a sensitive method for reviewing the performance of shielding configurations.
Figure 6: Steady-state sample temperature comparison across shielding configurations: a) Internal heating case with the superconducting quantum circuit itself as the heat source; b) External heating case with an absorptive low-pass filter positioned near the sample mount as the heat source.
Our results demonstrate that the sample temperature is most strongly influenced by the nearest surface – the sample holder lid. This conclusion establishes the sample holder itself as a critical component of the shielding system that must be optimized alongside traditional radiation shields. Inner coating of the sample mount should be considered optimal. It provides efficient heat dissipation path for SQC which is crucial for "floating" qubits where the qubit capacitor pad is electrically isolated. Outer coating on a single absorptive-coated shield around the mount is also optimal. It ensures effective thermal management of the parasitic heat sources, e.g. control wiring and filters. However, the simulation reveals a performance plateau starting from the "E" configuration, where additional shielding layers (configurations "G", and "H") do not provide any improvements. While basic infrared shielding is essential, multilayer shielding leads to unnecessary design complexity that increases system mass and footprint without improving thermal isolation.
Therefore, the following recommendations for shielding configurations are proposed:
  1. 1.
    Floating architecture qubits with nearby IR filters:
    Configuration “G” (inner-coated sample holder + single absorptive shield) provides optimal thermal isolation from both intrinsic qubit heating and nearby filter radiation.
  2. 2.
    Floating architecture without IR filters in a proximity:
    Configuration “C” (inner-coated sample holder) offers sufficient protection while minimizing system complexity.
  3. 3.
    Grounded qubit architectures with adjacent passive components:
    Configuration “F” (uncoated holder + single absorptive shield) effectively protects from parasitic heat loads from control wiring and other components.
  4. 4.
    Minimal requirements (no floating elements or warm components):
    Configuration “A” (basic uncoated copper holder) provides adequate baseline shielding without unnecessary mass or complexity.
Optimal SQC shielding requires careful geometry selection alongside absorber placement. Our analysis shows that about 35% of designs use light-tightHazard (2019) cup-lid configurationsPlace et al. (2021); Zhang (2018); Grünhaupt et al. (2018); Reagor (2015), often with indium-sealed or jointsHarrington (2020). Cylindrical inserts are less common Byoung-moo (2021); Chou (2018); Serniak (2019); Wang et al. (2019). Cup-lid configuration can be easily integrated in standard dilution refrigerators. Hermetic RF and DC connectors can be used to allow control wiring inside such a shielding. However, a vacuum pumping path with IR-tight design should be introduce to ensure efficient evacuation.
@@
In order to ensure a robust thermal contact, approximately in a half of considered papers, oxygen-free copper Buks et al. (2020); Eddins et al. (2019); Harrington (2020) is used (see Figure 7). The choice of a copper is justified by its high thermal conductivity. The holders manufactured of superconductors, such as aluminum Jin et al. (2015); Burkhart et al. (2021); Ribeill (2016); Kelly (2015), considered in the other half of papers. The use of superconductors creates additional magnetic shielding owing to the Meissner effect (more detail consideration of this effect will be done later). Despite the added protection from the magnetic field, superconducting sample holders have a lower thermal conductivity compared to the copper sample holders, that is why the use of the superconductors is undesirable. There are also copper sample holders coated by superconductor (Al Place et al. (2021) or InNarla (2017)). These sample holders create additional magnetic shielding while maintaining the high thermal conductivity. However, these examples are not common. There are also a few cases, where gold-coated copper is used Lienhard et al. (2019); Blok et al. (2021). The gold coating prevents the formation of the oxides on the copper, thereby reducing dielectric losses Huang et al. (2021).
@@
Among studies documenting sample holder geometry, more than 50% feature transmon qubits mounted within three-dimensional cavities Monroe et al. (2021b); Serniak (2019); Smith (2019); Grünhaupt et al. (2018); Martin et al. (2020); Chapman et al. (2023) that simultaneously function as both chip holders and readout circuit elements. The remaining implementations employ three distinct holder configurations: approximately 9% utilize rectangular holdersMaleeva et al. (2018); Ribeill (2016), 3% incorporate cylindrical holdersChen (2018); Quintana (2018), and 27% employ flat copper bases with printed circuit boards and protective lids positioned above the qubit chip.Byoung-moo (2021); Arute et al. (2019); Zhang (2018).
The selection of a sample holder’s shape depends on the SQC design and the wiring method to the cryogenic setup. Only the quality (roughness and absorption coefficient) of the sample holder internal surface and its sealing are of importance.
In 8% of the surveyed studies, the sample holder was hermetically sealed using indium wire to achieve light-tight conditionsChou (2018); Schwartz (2016); Narla (2017), while only 1.6% incorporated lids with IR-absorbent coatings Kreikebaum et al. (2016). For the latter case, the coating consisted of Stycast epoxy mixed with carbon lampblack and 350 ${\mu}m$ SiC particles. These experimental implementations show good agreement with our prior theoretical calculations.
The absorbing coating on the inner surface of the sample holder allows for more efficient cooling of the SQC, but only when the holder itself has high thermal conductivity. However, high dielectric loss of the absorber can affect qubit‘s coherence if they are close to each other. IR absorber should be positioned sufficiently far from the qubit. Sufficient distance can be determined using surface participation method Lienhard et al. (2019); Smirnov et al. (2024). This approach shows that the minimal qubit-absorber distance should exceed 4 $mm$. The surface participation method allows to quantify the extent to which the absorber interacts with the qubit’s field in terms of maximal possible coherence time $T_{1}$ that can be achieved in presence of the dielectric loss channel:
@@
where $T_{1}$, $\omega$, and $Q$ are coherence time, angular frequency and quality factor, respectively. $tan{\delta}_{abs}$ is the dielectric loss tangent of the absorber, $p_{abs}$ is a participation ratio of the material (a fraction of electromagnetic field energy stored within this material). ${\Gamma}_{0}$ represents other loss channels. The simulation conducted with a typically sized floating transmon Smirnov et al. (2024) and a usual transmon on a lossless substrate in a lossless chip cavity to ensure ${\Gamma}_{0}=0$. Inner side of the top lid was covered with RF absorber with thickness of 0.5 $mm$ and $tan{\delta}_{abs}$ of 0.05, which is an upper bound for Stycast IR absorber. During the simulation distance between the qubit and the bottom surface of the absorber was varied from 0.5 to 5 $mm$. The modelling shows that the influence of the absorber becomes as low as an influence of dielectric losses in silicon (limiting $T_{1}$ more than 10 $ms$) substrate when Stycast is more than 4 $mm$ away from the qubit (see Figure 8). This ensures that the benefits of efficient cooling provided by the absorber are not compromised by degradation in qubit coherence.
@@
Based on comprehensive radiation modeling and literature review, we establish three critical design requirements for optimal sample holder performance:
  1. 1.
    Copper (or copper with superconducting coating) provides optimal thermal conductivity for efficient SQC cooling, balancing both heat dissipation and magnetic shielding requirements.
  2. 2.
    Indium wire gaskets offer reliable hermetic sealing of the holder lid, preventing photon leakage while maintaining cryogenic compatibility down to millikelvin temperatures.
  3. 3.
    The inner surfaces of holders for floating-architecture SQCs require absorptive coatings to mitigate infrared radiation while preserving quantum coherence.
@@ II.2.1 Shielding principle
Generally, protection from electrical and magnetic fields consists of three components: electrostatic shielding, magnetostatic shielding, and electromagnetic shielding (Figure 9). The principle of shielding from an electrical field (Figure 9a) involves the accumulation of the charge from a free space on the shield and its subsequent leakage into the ground. For effective protection against an electrical field, the shield must be made of a highly conductive material, such as copper or aluminum, with high-quality electrical grounding to minimize the contact resistance with the ground.
The shields made of ferromagnetic materials, such as permalloys (${\mu}$-metal, a soft magnetic alloy, or steels), with a high relative magnetic permeability (${\mu}_{r}$) are used to protect from a static and slowly varying (up to 1 $kHz$) magnetic field. In these shields, the lines of magnetic flux pass along the walls (see Figure 9b), which have a lower magnetic resistance than the surrounding air. The shielding effectiveness is determined by magnetic permeability of the material the shield is made of.
The operation principle of the shields protecting from an alternating high-frequency magnetic field is based on an alternating electromagnetic field excitation within the shield. It creates alternating induction eddy currents (also known as Foucault currents). These currents shield the volume which this currents surround: the field inside the shield is brought down, while the field outside the shield is enhanced (see Figure 9c). As a result, the field strength is reduced inside the shield and increased outside, thus leading to the field to be pushed from the shield. This type of shielding depends on the depth of field penetration at various frequencies (known as the "skin depth"), and becomes effective above frequencies ($f$) 1 $kHz$. The lower the skin depth ($\delta$) the thinner (the lighter) the shield can be made:
@@ II.2.3 Practical implementation of electromagnetic shielding
Superconducting shields used for protection from parasitic magnetic fields are primarily manufactured from aluminum (Figure 10) Jin et al. (2015); Moghaddam et al. (2019); Asaad et al. (2016); Liu (2016); Singh et al. (2025); Warren et al. (2023); Norris et al. (2024). Additional superconducting materials identified in the literature include tin (Sn)Kreikebaum et al. (2016) and indium (In)Premkumar et al. (2021). Aluminum is preferred material for superconducting shielding due to the ease of its workability and relatively low cost compared to other materials.
In considered papers, almost a quarter of all superconducting shields in combination with an IR shielding, which are coated with an IR-absorbing material, are used Zhang (2018); Arute et al. (2019); Kreikebaum et al. (2016). These shields usually consist of a light tight cup with a lid Place et al. (2021); Zhang (2018), which allows for more compact shielding due to the possibility of combining together IR and magnetic shielding (shielding from magnetic field). However, superconductors have a lower thermal conductivity than in its normal state. It has a negative impact on the absorbed heat dissipation, which is crucial for IR shielding. To effectively cool a superconducting shield, a superconducting layer can be placed on a copper base Premkumar et al. (2021); Kreikebaum et al. (2016).
@@
The operation principle of magnetic shielding is based on the magnetic field retraction into a shield material that has a high relative magnetic permeability (typically 100,000 or more) Huang et al. (2021). These components are typically fabricated from permalloy (iron-nickel alloys) that undergo specialized processing, including hydrogen atmosphere annealing. In the literature, these materials are often referred to as ${\mu}$-metals Hazard (2019); Jr. (2018); Andersen et al. (2019); Koolstra et al. (2022); Wilen et al. (2021) or they may have specific names (Figure 11), such as Cryoperm Lisenfeld et al. (2019); Sung (2018); Ding et al. (2023); Dodge et al. (2023), Cryophy Dickel et al. (2018); Asaad et al. (2016); Norris et al. (2024), Amumetal Wang (2022); Narla (2017); Liu (2016); Béjanin et al. (2016).
The presented materials differ in the method of their thermal processing in a hydrogen environment, and consequently, in their relative magnetic permeability (Table 2). For these shields, it is preferable to use a material with the highest relative magnetic permeability.
For magnetic shielding applications, cup-shaped shield configurations are commonly used Bengtsson (2020); Norris et al. (2024). Similarly, cylindrical shields with lids demonstrate comparable prevalence in the literature Béjanin et al. (2016); Quintana (2018); Zhang (2018). Individual magnetic shielding is used in special cases Arute et al. (2019) As with superconducting shielding, the presence of a lid on the shield enables the equalization of the magnetic field within the protected space, thus it is recommended to use the magnetic shield with a lid.
@@
In some studies, specific guidelines are provided for the use of magnetic shields. For example, it is recommended not to include magnetic components inside the shield Janzen et al. (2023); Reagor (2015); Opremcak et al. (2021); Gao (2018), or the sample holder should be placed closer to the bottom of the shield Underwood (2015), and a thermal anchor should be installed on the shield to improve its cooling Zhang (2018); Reagor (2015); Place et al. (2021). Some magnetic shields have special design features such as waveguides on the lid Byoung-moo (2021) or light tightness Hazard (2019). These design features help to minimize the influence of the magnetic field on the SQC. Almost a half of considered shielding systems have the sample holders with special features. Either the holder itself is made of a superconductor Chen (2018); Mundhada (2020); Wilen et al. (2021), or it has superconducting Place et al. (2021) or absorbing layers Kreikebaum et al. (2016). These aspects have been discussed in more details before in a section about sample holders.
Figure 12: Simulation of magnetic field distribution within cylindrical shields: a) calculation model (the external magnetic field of 50 $\mu$T directed along the axis of the cylinders), violet dashed line is modeling area and green area is average value area calculation for magnetic field suppression (MFS) determination; b) comparison of magnetic field strength along the shield axis for various shielding configurations; magnetic field distribution within: c) single $\mu$-metal shield (color bar for magnetic field strength is also valid for d)-g) configurations; the boundaries of the simulated shields are indicated by a black dashed line here and on the d)-g) various configurations; the MFS is indicated in dB on the lower third part of the shielding system here and for d)-g) configurations); d) single aluminum shield, e) double $\mu$-metal shield, f) aluminum shield surrounded with a $\mu$-metal shield, g) $\mu$-metal shield surrounded with an aluminum shield.